OpenStudy (anonymous):

polynomial difference of (2x^2+6x-7)-(5x^2+3x) equals what

hero (hero):

(2x^2+6x-7)-(5x^2+3x) = 2x^2 + 6x - 7 - 5x^2 -3x = 2x^2 - 5x^2 + 6x - 3x - 7 = -3x^2 + 3x - 7